OpenStudy (anonymous):

Will give medal + fan. In 3 hours, Jim can walk 14 miles. To find the time he would take to walk 25 miles, he wrote the proportion 3/14 = 25/h. Explain his error. Please help! Thanks :)

OpenStudy (cj49):

3 14 x 25 25x3/14=?

OpenStudy (swissgirl):

Ok so basically its all abt propotions its hours/miles walked so for 3 hours he walked 14 miles then how many hours will it take him to walk 25 miles \[\frac{3 hours}{14 miles}=\frac{x hours}{25 miles}\]

OpenStudy (swissgirl):

They set it up incorrectly and confused the hours with the miles

OpenStudy (anonymous):

Okay. Thanks so much @swissgirl :)